OpenStudy (anonymous):

if f(x)=x-10 and g(x)=4x+3, how do i find fog of (-2)

OpenStudy (anonymous):

fog (-2)=f( g(-2))=f(4(-2)+3)=f(-5)=-5-10=-15

OpenStudy (tyteen4a03):

What is "fog"?

OpenStudy (anonymous):

the composition of f and g

OpenStudy (anonymous):

First you need to take the function g(x) and substitute the entire thing for x in f(x). The resulting function would be f(g(x)) then the next thing is simply to substitute -2 for x in f(g(x))